OpenStudy (anonymous):

Yulian works at the zoo feeding the animals. He puts water in the elephant habitat at the beginning of the day. The table shows how much water, in gallons, remains at different times during the day. What is the equation that represents Yulian’s situation? (y – 29) = ___(x –___ ) Yulian’s Situation Hours 2 4 6 8 Water remaining 43 36 29 22

OpenStudy (bibby):

point slope form equations are of the form \(y-y_1=m(x-x_1)\) where \((x_1,y_1)\) is a point the line goes through and m is the slope

OpenStudy (anonymous):

i still don't understand... @bibby

OpenStudy (bibby):

sorry, the site has been acting up all day for me (x1,y1) is one of the points the line or function goes through for example a line with a slope of 5 that passes through (3,-6) would look like \(y--6=5(x-3)\) \(y+6=5(x-3)\)

OpenStudy (anonymous):

so that's the answer?

OpenStudy (bibby):

no? I don't give answers, I give you the recipe

OpenStudy (anonymous):

oh.. hahaha thanks

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

it's easy to give you the answer it's more rewarding to teach you /how/

OpenStudy (anonymous):

yeah.. i know.. but i'm in the middle of a test and was hoping for a quick answer. @bibby

OpenStudy (bibby):

we can't help with tests. the only help I can give is that slope = \(m=\large \frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

im just not good at math... @bibby

OpenStudy (bibby):

I suck too. I still try sometimes

OpenStudy (anonymous):

can i please have the answer? @bibby

OpenStudy (bibby):

ask nicely wait damn you said the magic word

OpenStudy (bibby):

you have 2 points (2,43) & (4,36) plug that into the formula and find the slope

OpenStudy (anonymous):

thx @bibby